OpenStudy (anonymous):

PLZ PLZ, SOMEONE HELP ME ON THIS PROBLEM!!! How many quarts of pure antifreeze should Kelly add to 10 quarts of a 20% antifreeze solution to make a 50% antifreeze solution?

OpenStudy (anonymous):

10 quarts is 20 percent, therefore you times by 5 to get 100 percent. So 100 percent = 50 quaarts and 50 percent = 25 quarts.

OpenStudy (anonymous):

can you please write that in number form..so i can see how it looks?

OpenStudy (anonymous):

20%=10 quarts 100%= 50 quarts 50%= 25 quarts

OpenStudy (anonymous):

get it?

OpenStudy (cwrw238):

or u can do it this way: let x = no. of quarts to be added then: 0.2* 10 + x = 0.5(x + 10) 2 + x = 0.5x + 5 o.5x = 3 x = 15 you need to add 15 quarts